# Variational Gaussian Approximation in Replica Analysis of Parametric Models

### v1

@@ I Introduction
The statistical mechanics analysis of learning and inference from examples has a long history, starting with the seminal work of Seung, Somplinsky and Tishby in the early 1990s Sompolinsky et al. (1990); Seung et al. (1992). Since then, such approaches have become a standard topic in the statistical mechanics of disordered systems Engel (2001); Nishimori (2001). One of the central analytical tools in this line of research is the replica method Mezard et al. (1986); Charbonneau et al. (2023); Montanari and Sen (2024). It provides a powerful method to evaluate the average behavior of statistical quantities such as free energy and prediction error. Unlike classical learning theory, which often yields only bounds for those quntities Shalev-Shwartz and Ben-David (2014), the replica method can give sharp predictions.
Despite its strength, the replica method has a major limitation. To take quenched averages analytically, one usually has to assume highly idealized data-generating distributions. This simplification is required even in problems as basic as the standard linear regression. Hence, the replica method has been regarded as a mathematical technique for guessing exact solutions under highly restricted conditions, rather than a tool for predicting universal structures of learning and inference, or for describing the behavior of them in realistic data, although some exceptions exist Malzahn and Opper (2000, 2001, 2002, 2003, 2005); Yoshino (2018); Loureiro et al. (2021); Yoshino (2020, 2023).
A particularly notable line of work in the context of learning and inference was provided by the series of studies by Malzahn and Opper Malzahn and Opper (2000, 2001, 2002, 2003, 2005). By combining the replica method with variational approximations, they successfully obtained an approximate solution to learning curves for Gaussian process regression and hard-margin support vector machines without assuming a concrete data distributions. The use of the variational approximation simultaneously offered two important advantages. They relax the dependence on overly simplified data distributions and, moreover, it allowed a systematic way of giving approximate predictions even when exact replica calculations were not trivial. These advantages may be useful in the analysis of learning with modern complex models.
@@ II Setup
The function $l(\theta;z_{i})$ represents a loss associated with each data point, such as a negative log-likelihood, and the term $\tfrac{\lambda}{2}\|\theta\|_{2}^{2}$ corresponds to the weight decay. The sum $\mathcal{L}$} corresponds to a Hamiltonian in statistical mechanics. More generally, regularization can be incorporated into the measure over $\theta$, which would be more suitable when the regularization is more complicated, or the parameter is discrete. The parameter $\beta>0$ is an inverse temperature. The case $\beta=1$ corresponds to Bayesian inference with the posterior distribution, while in the limit $\beta\to\infty$ the distribution concentrates on the minimizers of ${\mathcal{L}}(\theta;D_{n})$, which corresponds to empirical risk minimization.
@@ III.1 Grand canonical formalism
$p_{r,{\rm GC}}^{\beta,n}(\theta^{1},\ldots,\theta^{r})$ $=\frac{1}{\Xi_{r,{\rm GC}}^{\beta,n}}e^{nH(\bm{\theta})},$ (13)
@@
From the viewpoint of statistical mechanics, this corresponds to the grand canonical ensemble where the number of data points, instead of the number of particles, fluctuates with chemical potential $\beta^{-1}\log n$. Hence, we refer to  (13) as the grand canonical replicated system. Appendix A presents a derivation more directly emphasizing the formal analogy with the grand canonical ensemble. When computing averages, either  (10) or  (13) can be used in principle, but as we shall see, the GC formalism often simplifies the analysis.
@@ III.2 General formulae for errors
In these expressions, the expectation with respect to the GC replicated system plays the role of a cavity bias at the data point $z$. In particular, the formula for the training error shows that the factor $e^{-\beta\sum_{a=1}^{r}l(\theta^{a};z)}$ appears as a bias because the evaluation point is included in the training dataset.
@@ IV Variational Gaussian Approximation (VGA)
where $\tilde{\mathbb{E}}_{r,{\rm GC}}^{\beta,n}[\cdot]$ denotes the average with respect to the approximate replicated system proportional to $e^{n\tilde{H}(\bm{\theta})}$. The derivation of this variational free energy follows the standard perturbative variational principle Parisi (1988) and, for integer $r$, it provides a bound on the true free energy. However, in the limit $r\to 0$, its leading term $\lim_{r\to 0}r^{-1}{\mathcal{F}}_{r,{\rm GC}}^{\beta,n}$ is not necessarily guaranteed to bound the true free energy $\mathbb{E}_{D_{n}}[\log Z^{\beta,n}(D_{n})]$. Hence, instead of naively maximizing or minimizing with respect to the parameters of the trial Hamiltonian, we have to consider stationary conditions. Appendix B illustrates a simple example in which the variational parameters are at a saddle point rather than a maximum or minimum. It should be noted that the stationarity conditions are considered solely from the requirement of choosing the trial Hamiltonian optimally. Hence, unlike in many derivations of exact solutions, taking the thermodynamic limit is not necessary here, although one could consider an appropriate limit later.
@@
It is worth emphasizing that the use of the grand canonical formulation yields the interaction term ${\mathcal{F}}_{\rm int}$ in a form where the expectation over the data distribution $z\sim p_{\rm{data}}$ is taken after averaging the factor $e^{-\beta\sum_{a=1}^{r}l(\theta^{a};z)}$ with respect to the approximate distribution. If we instead use the original formulation  (10), one would have to deal with the term $\log\mathbb{E}_{z\sim p_{\rm{data}}}[e^{-\beta\sum_{a=1}^{r}l(\theta^{a};z)}]$ inside the average with respect to the approximate distribution. Such a factor is difficult to handle unless the expectation over $p_{\rm{data}}$ can be computed analytically. This makes progress almost impossible unless the average over $p_{\rm{data}}$ is analytically tractable. By contrast, in the present formulation one may obtain a tractable expression for the interaction term by averaging with respect to the approximate distribution, provided that the trial Hamiltonian is sufficiently simple.
In practice, we derive the stationarity conditions while keeping the expectation over $p_{\rm{data}}$ explicit, and when needed we replace it by the empirical average. This makes it possible to obtain approximate formulae that apply to general datasets.
@@ V.1 Stationarity conditions
These conditions correspond, respectively, to the stationarity of the population loss gradient, a (non-centered) covariance-like quantity for the gradients, and a regularized Hessian-like quantity. They are, however, evaluated under thermal averages, and therefore do not coincide with naive population quantities at finite $n$.
@@ VI.1 General Properties
In practice, even when the data-generating distribution $p_{\rm{data}}$ is unknown, the expectations over $z$ that appear in the above expressions can be approximated by empirical averages over the observed data. This allows the generalization gap to be estimated for arbitrary distributions.
@@ VI.2 Linear regression
$\bar{\epsilon}_{\rm tr}$ $=\mathbb{E}_{z}\left[\frac{1}{(1+\chi_{f}(x))^{2}}\,(y-m_{f}(x)-\xi_{f})^{2}\right],$ (73)
$\bar{\epsilon}_{\rm pred}$ $=\mathbb{E}_{z}\left[(y-m_{f}(x)-\xi_{f})^{2}\right],$ (74)
$\bar{\delta\epsilon}$ $=\mathbb{E}_{z}\left[\Bigl{(}1-\frac{1}{(1+\chi_{f}(x))^{2}}\Bigr{)}(y-m_{f}(x)-\xi_{f})^{2}\right].$ (75)
@@ VI.3 Application to concrete problems
The most notable feature of the present formulation is that the parameters of the trial Hamiltonian can be determined adaptively for a given data-generating distribution. By approximating the population average $\mathbb{E}_{z\sim p_{\rm{data}}}$ with the empirical distribution of the observed data, the method can be applied not only to synthetic data with analytically tractable expectations, but also to real datasets where the true distribution is unknown.
In this subsection, we illustrate this point with several concrete settings. For each case, we approximate $\mathbb{E}_{z}$ by the empirical average over the observed data to determine $m_{w},q_{w},\chi_{w}$, and then evaluate the predicted generalization error $\bar{\epsilon}_{\rm pred}$. It is worth emphasizing that even with moderately large datasets, quantifying the fluctuations represented by $\chi_{w}$ and $q_{w}$ is not at all trivial without repeated experiments, yet the VGA provides a principled framework to estimate them.
@@
(b) 2NN (relu), $\lambda=10^{-2}$
(c) 2NN (tanh), $\lambda=10^{-2}$
(d) YP MSD, $\lambda=10^{-2}$
(e) linear, $\lambda=10^{-4}$
@@
(g) 2NN (tanh), $\lambda=10^{-4}$
(h) YP MSD, $\lambda=10^{-4}$
Figure 1: Verification of variational Gaussian approximation (VGA) in linear regression. Markers represent true generalization errors evaluated by large test data. Lines are predictions of VGA.
@@ VI.3.1 Synthetic data: teacher–student scenario
Based on this generative model, we construct datasets of size $10^{4}$ and use them to approximate the averages over $z\sim p_{\rm{data}}$, evaluate the predicted generalization error, and compare with simulations. Note that this size $10^{4}$ is different from $n$ that appears in theoretical analysis, which is merely a control parameter. Since the components of $x$ are i.i.d., we restrict ourselves to the simplified structure $q_{w}=\bar{q}_{w}I_{d}$ and $\chi_{w}=\bar{\chi}_{w}I_{d}$.
Figure 1 (a)–(c), (e)–(g) show the comparison between the predictions of VGA and the experimental results. In all cases, the variational approximation provides a quantitatively accurate description.
@@ VI.3.2 Real-world data
We next consider real-world data, using the Year Prediction MSD (YPMSD) dataset Bertin-Mahieux (2011). This dataset consists of the task of predicting the release year of a song from 90-dimensional audio features, with a total of 515,345515{,}345 samples.
In our setting, we assume that only 20% of the data are available, on which the VGA is performed, while the remaining 80% are hidden and used as a test set to estimate the true generalization error. All features are standardized to have zero mean and unit variance.
For the variational approximation, we consider two settings for $\chi_{w}$ and $q_{w}$: one restricted to diagonal matrices and the other allowing full matrices. The results are shown in Figure 1(d) and (h). In both cases the variational predictions are consistent with the empirical generalization errors when the full matrices are used, while the diagonal approximation exhibits deviations for small sample sizes. This indicates that correlations among features play a crucial role in this dataset.
These results demonstrate that the validity of the variational approximation extends beyond controlled synthetic settings to real-world datasets
@@ VII Summary and Conclusion
To this end, we introduced a grand canonical formalism that replaces dataset averages by virtual sampling from an infinitely large data reservoir, and applied a variational approximation to the resulting replicated system. Within this framework, the stationarity conditions for the variational parameters can be derived without performing an analytic average over the data, allowing them to be determined adaptively for each given data distribution. A key technical point of our approach is that it changes the standard assumption in replica analyses: instead of requiring that the average over data can be carried out analytically, our method only requires that the trial Hamiltonian is sufficiently simple. This perspective revealed connections to well-known information criteria in statistics and machine learning. As a concrete application, we analyzed linear regression and demonstrated that the method can yield learning curves even for problems involving real-world datasets, where the data-generating distribution is not known explicitly. Although the basic idea of GC formalism was already given by Malzahn and Opper Malzahn and Opper (2002, 2000, 2001, 2005), we believe the present work provides a clearer view for analysis in parameter space.
Natural extensions of this work include analyses of more complex models, such as multilayer neural networks, for which exact solutions have not been obtained yet except for limited scenarios. Although the analysis of neural networks can become highly involved even with simple trial Hamiltonians, promising directions include structured scenarios such as sparse teacher–student models Yoshino (2020, 2023). Also, in recent years, variational approximation methods with neural networks have advanced significantly. In this work we focused on the classical variational Gaussian approximation for the sake of analytical tractability, but the development of numerical variational methods would be also an important direction for future research.

### v2

@@ I Introduction
The statistical mechanics analysis of learning and inference from examples has a long history, starting with the seminal work of Sompolinsky, Tishby, and Seung in the early 1990s Sompolinsky et al. (1990); Seung et al. (1992). Since then, such approaches have become a standard topic in the statistical mechanics of disordered systems Engel (2001); Nishimori (2001). One of the central analytical tools in this line of research is the replica method Mezard et al. (1986); Charbonneau et al. (2023); Montanari and Sen (2024). It provides a powerful method to evaluate the average behavior of statistical quantities such as free energy and prediction error. Unlike classical learning theory, which often yields only bounds for those quntities Shalev-Shwartz and Ben-David (2014), the replica method can give sharp predictions.
Despite its strength, the replica method has a major limitation. To take quenched averages analytically, one usually has to assume highly idealized data-generating distributions. This simplification is required even in problems as basic as the standard linear regression. Hence, the replica method has been regarded as a mathematical technique for guessing exact solutions under highly restricted conditions, rather than a tool for predicting universal structures of learning and inference, or for describing the behavior of them in realistic data, although some exceptions exist Malzahn and Opper (2000, 2001, 2002, 2003a, 2005); Yoshino (2018); Loureiro et al. (2021); Yoshino (2020, 2023).
A particularly notable line of work in the context of learning and inference was provided by the series of studies by Malzahn and Opper Malzahn and Opper (2000, 2001, 2002, 2003a, 2005). By combining the replica method with variational approximations, they successfully obtained an approximate solution to learning curves for Gaussian process regression and hard-margin support vector machines without assuming a concrete data distributions. The use of the variational approximation simultaneously offered three important advantages. They relax the dependence on overly simplified data distributions, and it allowed a systematic way of giving approximate predictions even when exact replica calculations were not trivial. Moreover, it does not necessarily require taking the thermodynamic limit. These advantages may be useful in the analysis of learning with modern complex models.
@@ II Setup
The function $l(\theta;z_{i})$ represents a loss associated with each data point, such as a negative log-likelihood, and the term $\tfrac{\lambda}{2}\|\theta\|_{2}^{2}$ corresponds to the weight decay. The sum $\mathcal{L}$} corresponds to a Hamiltonian in statistical mechanics. More generally, regularization can be incorporated into the measure over $\theta$, which would be more suitable when the regularization is more complicated, or the parameter is discrete. The parameter $\beta>0$ is an inverse temperature. The case $\beta=1$ corresponds to the generalized Bayesian inference based on the score $l(\theta;z_{i})$, which reduces to ordinary Bayesian inference when $l$ is the negative log-likelihood. On the other hand, in the limit $\beta\to\infty$ the distribution concentrates on the minimizers of ${\mathcal{L}}(\theta;D_{n})$, which corresponds to empirical risk minimization.
@@ III.1 Grand canonical formalism
$p_{r,{\rm GC}}^{\beta,n}(\bm{\theta})$ $=\frac{1}{\Xi_{r,{\rm GC}}^{\beta,n}}e^{nH(\bm{\theta})},$ (13)
@@
In fact, this rewriting is similar to the transform from canonical ensemble to the grand canonical ensemble in statistical mechanics where the number of data points, instead of the number of particles, fluctuates. Hence, we refer to (13) as the grand canonical replicated system. See Appendix A for a derivation more directly emphasizing the formal analogy with the grand canonical ensemble. When computing averages, either (10) or (13) can be used in principle, but as we shall see, the GC formalism often simplifies the analysis.
@@ III.2 General formulae for errors
In these expressions, the expectation with respect to the GC replicated system plays the role of a cavity bias at the data point $z$. In particular, the formula for the training error shows that the factor $e^{-\beta\sum_{a=1}^{r}l(\theta^{a};z)}$ appears as a bias due to the inclusion of the evaluation point in the training dataset.
@@ IV Variational Gaussian Approximation (VGA)
where $\tilde{\mathbb{E}}_{r,{\rm GC}}^{\beta,n}[\cdot]$ denotes the average with respect to the approximate replicated system $e^{n\tilde{H}(\bm{\theta})}/\tilde{\Xi}_{r,{\rm GC}}^{\beta,n}$, where $\tilde{\Xi}_{r,{\rm GC}}^{\beta,n}$ is the normalization constant. The derivation of this variational free energy follows the standard perturbative variational principle Parisi (1988); Opper and Saad (2001); Murphy (2023) and, for integer $r$, it provides a bound on the true free energy. However, in the limit $r\to 0$, its leading term $\lim_{r\to 0}r^{-1}{\mathcal{F}}_{r,{\rm GC}}^{\beta,n}$ is not necessarily guaranteed to bound the true free energy $\mathbb{E}_{D_{n}}[\log Z^{\beta,n}(D_{n})]$. Hence, instead of naively maximizing or minimizing with respect to the parameters of the trial Hamiltonian, we have to consider stationary conditions. Appendix B illustrates a simple example in which the variational parameters are at a saddle point rather than a maximum or minimum. It should be noted that the stationarity conditions are considered solely from the requirement of choosing the trial Hamiltonian optimally. Hence, unlike in many derivations of exact solutions, taking the thermodynamic limit is not necessary here, although one could consider an appropriate limit later.
@@
It is worth emphasizing that the use of the GC formulation yields the interaction term ${\mathcal{F}}_{\rm int}$ in a form where the expectation over the data distribution $z\sim p_{\rm{data}}$ is taken after averaging the factor $e^{-\beta\sum_{a=1}^{r}l(\theta^{a};z)}$ with respect to the approximate distribution. If we instead use the original formulation (10), one would have to deal with the term $\log\mathbb{E}_{z\sim p_{\rm{data}}}[e^{-\beta\sum_{a=1}^{r}l(\theta^{a};z)}]$ inside the average with respect to the approximate distribution. Such a factor is difficult to handle unless the expectation over $p_{\rm{data}}$ can be computed analytically. This makes progress almost impossible unless the average over $p_{\rm{data}}$ is analytically tractable. By contrast, in the present formulation one may obtain a tractable expression for the interaction term by averaging with respect to the approximate distribution, provided that the trial Hamiltonian is sufficiently simple.
In practice, we derive the stationarity conditions while keeping the expectation over $p_{\rm{data}}$ explicit, and when needed we approximate it with the empirical average of the observed data. This makes it possible to obtain approximate formulae that apply to general datasets.
@@ V.1 Stationarity conditions
These conditions correspond, respectively, to the stationarity of the population loss gradient, a (non-centered) covariance-like quantity for the gradients, and a regularized Hessian-like quantity. They are, however, evaluated under thermal averages, and therefore do not coincide with naive population quantities when $n$ is finite.
@@ VI.1 General Properties
In practice, even when the data-generating distribution $p_{\rm{data}}$ is unknown, the expectations over $z$ that appear in the above expressions can be approximated by empirical averages over the obtained data. This allows the generalization gap to be estimated for arbitrary distributions.
@@ VI.2 Linear regression
$\bar{\epsilon}_{\rm tr}$ $=\mathbb{E}_{z}\left[\frac{1}{(1+\chi_{f}(x))^{2}}\bigl((y-m_{f}(x))^{2}+q_{f}(x)\bigr)\right],$ (73)
$\bar{\epsilon}_{\rm pred}$ $=\mathbb{E}_{z}\left[(y-m_{f}(x))^{2}+q_{f}(x)\right],$ (74)
$\bar{\delta\epsilon}$ $=\mathbb{E}_{z}\left[\Bigl(1-\frac{1}{(1+\chi_{f}(x))^{2}}\Bigr)\bigl((y-m_{f}(x))^{2}+q_{f}(x)\bigr)\right].$ (75)
@@ VI.3 Application to concrete problems
The most notable feature of the present formulation is that the parameters of the trial Hamiltonian can be determined adaptively for a given data-generating distribution. By approximating the population average $\mathbb{E}_{z\sim p_{\rm{data}}}$ using some data, the method can be applied not only to synthetic data with analytically tractable expectations, but also to real datasets where the true distribution is unknown. In this subsection, we illustrate this point with several concrete settings.
For each case, we approximate the average $\mathbb{E}_{z}[{\mathcal{O}}(z)]$ of a statistic ${\mathcal{O}}(z)$ by the empirical average over the data of size $n_{0}$ as $n_{0}^{-1}\sum_{i=1}^{n_{0}}{\mathcal{O}}(z_{i})$. Using this approximation with fixed $n_{0}$, we numerically solve the stationarity conditions (70)-(72) to determine $m_{w},q_{w},\chi_{w}$, and then predict the generalization error $\bar{\epsilon}_{\rm pred}$. That is, we approximate $\mathbb{E}_{z}$ using a dataset of size $n_{0}$, and predict the learning behavior when trained on data of size $n$. It should be noted that, in general, $n_{0}$ is independent of $n$. Given $_{0}, nn$ is merely a control parameter. Thus, both $n\leq n_{0}$ and $n\geq n_{0}$ are possible. Of course, the accuracy of the final result depends on $n_{0}$, as discussed in subsubsection VI.3.2.
We also remark that even with moderately large datasets, quantifying the fluctuations represented by $\chi_{w}$ and $q_{w}$ is not at all trivial without repeated experiments, yet the VGA provides a principled framework to estimate them.
@@
(b) linear, $\lambda=10^{-4}$
(c) linear, $\lambda=10^{-2}$
(d) linear, $\lambda=10^{-4}$
(e) 2NN (relu), $\lambda=10^{-2}$
@@
(g) 2NN (relu), $\lambda=10^{-2}$
(h) 2NN (relu), $\lambda=10^{-4}$
(i) 2NN (tanh), $\lambda=10^{-2}$
(j) 2NN (tanh), $\lambda=10^{-4}$
(k) 2NN (tanh), $\lambda=10^{-4}$
(l) 2NN (tanh), $\lambda=10^{-2}$
(m) YP MSD, $\lambda=10^{-2}$
(n) YP MSD, $\lambda=10^{-4}$
(o) YP MSD, $\lambda=10^{-2}$
(p) YP MSD, $\lambda=10^{-4}$
Figure 1: Verification of variational Gaussian approximation (VGA) in linear regression. Both generalization and training errors are shown. (a)-(l): synthetic data. (m)-(p): real-world data.Markers represent true generalization and training errors evaluated by large data of size $n_{0}=10^{4}$. Lines are predictions of VGA.
@@ VI.3.1 Synthetic data: teacher–student scenario
Based on this generative model, we construct datasets of size $n_{0}=10^{4}$ and use them to approximate the averages over $z\sim p_{\rm{data}}$, evaluate the predicted generalization error, and compare with simulations. Recall that this size $n_{0}$ can be different from $n$ that appears in stationarity condition.
Since the components of $x$ are i.i.d., we restrict ourselves to the simplified structure $q_{w}=\bar{q}_{w}I_{d}$ and $\chi_{w}=\bar{\chi}_{w}I_{d}$.
Figure 1 (a)–(l) show the comparison between the predictions of VGA and the experiments. In all cases, VGA provides a quantitatively accurate prediction.
@@ VI.3.2 Real-world data
(a) $(n_{0},\lambda)=(515,10^{-2})$
(b) $(n_{0},\lambda)=(1031,10^{-2})$
(c) $(n_{0},\lambda)=(2061,10^{-2})$
(d) $(n_{0},\lambda)=(5153,10^{-2})$
(e) $(n_{0},\lambda)=(515,10^{-4})$
(f) $(n_{0},\lambda)=(1031,10^{-4})$
(g) $(n_{0},\lambda)=(2061,10^{-4})$
(h) $(n_{0},\lambda)=(5153,10^{-4})$
(i) $(n_{0},\lambda)=(515,10^{-2})$
(j) $(n_{0},\lambda)=(1031,10^{-2})$
(k) $(n_{0},\lambda)=(2061,10^{-2})$
(l) $(n_{0},\lambda)=(5153,10^{-2})$
(m) $(n_{0},\lambda)=(515,10^{-4})$
(n) $(n_{0},\lambda)=(1031,10^{-4})$
(o) $(n_{0},\lambda)=(2061,10^{-4})$
(p) $(n_{0},\lambda)=(5153,10^{-4})$
Figure 2: The dependence of Learning curves for the YP MSD dataset on $n_{0}$. (a)-(h): generalization error $\bar{\epsilon}_{\mathrm{pred}}$. (i)-(p): training error $\bar{\epsilon}_{\rm tr}$. The average $\mathbb{E}_{z\sim p_{\rm{data}}}$ is approximated using a dataset of size $n_{0}$. Blue lines: VGA predictions. Red circles: experiments. Cyan dashed line: $n_{0}/d$.
We next consider real-world data, using the Year Prediction MSD (YPMSD) dataset Bertin-Mahieux (2011). This dataset consists of the task of predicting the release year of a song from 90-dimensional audio features, with a total of 515,345515{,}345 samples. All features are standardized to have zero mean and unit variance.
For the choice of $n_{0}$, we first use 20% of the data to approximate $\mathbb{E}_{z}$ ($n_{0}\simeq 10^{5}$), and the remaining 80% to estimate the true generalization error. Later, the dependence of the predictions on $n_{0}$ will be discussed. For the structure of $\chi_{w}$ and $q_{w}$, we consider two settings: one restricted to diagonal matrices and the other allowing full matrices.
The results for $n_{0}\simeq 10^{5}$ are shown in Figure 1(m)-(p). In all cases VGA yields quantitatively accurate predictions when full matrices are used, while the diagonal approximation exhibits deviations for small sample sizes. This indicates that correlations among features play a crucial role in this dataset. These results demonstrate that the validity of the variational approximation extends beyond controlled synthetic settings to real-world datasets.
Next, we check the dependence of the predictions on $n_{0}$. Figure 2 illustrates the dependence of the learning curves on $n_{0}$. Here, we varied $n_{0}$ from $515$ to $5153$, and for each estimate of the learning result with data size $n$, a different subset of size $n_{0}$ was sampled from the overall dataset to approximate the average over $z$. Also, for $q_{w}$ and $\chi_{w}$, only full matrices are considered. For $n_{0}\simeq 500$, the estimates deviate from the true values and have large fluctuations, while for $n_{0}\simeq 2000$, the estimates become almost exact. It should be noted that accurate predictions are obtained even for $n\gg n_{0}$. In this sense, our approach is essentially different from bootstrap analyses Efron (1979); Efron and Tibshirani (1994), where the variance vanishes when extrapolation to large data sizes is considered, even though the core idea behind the derivation of the GC formalism is similar to the bootstrap method.
@@ VII Summary and Conclusion
To this end, we introduced a grand canonical formalism that replaces dataset averages by virtual sampling from an infinitely large data reservoir, and applied a variational approximation to the resulting replicated system. Within this framework, the stationarity conditions for the variational parameters was derived without performing an analytic average over the data, allowing them to be determined adaptively for each given data distribution. A key technical point of our approach is that it changes the standard assumption in replica analyses: instead of requiring that the average over data can be carried out analytically, our method only requires that the trial Hamiltonian is sufficiently simple. This perspective revealed connections to well-known information criteria in statistics and machine learning. As a concrete application, we analyzed linear regression and demonstrated that the method can yield learning curves even for problems involving real-world datasets, where the data-generating distribution is not known explicitly. Although the basic idea of GC formalism was already given by Malzahn and Opper Malzahn and Opper (2002, 2000, 2001, 2005), we believe the present work provides a clearer view for analysis in parameter space.
Natural extensions of this work include analyses of more complex models, such as multilayer neural networks, for which exact solutions have not been obtained yet except for limited scenarios. Although the analysis of neural networks would become highly involved even with simple trial Hamiltonians, promising directions include structured scenarios such as sparse teacher–student models Yoshino (2020, 2023). Also, in recent years, non-parametric variational approximation methods with neural networks have advanced significantly. Hence, the development of numerical variational methods would be also an important direction. Furthermore, hypothesis testing methods based on replica analysis have often relied on state evolution equations of solvable settings Javanmard and Montanari (2014); Takahashi and Kabashima (2018); Sur and Candès (2019); Na et al. (2023), and our approach may relate to directions aiming to propose methods with a more direct connection to observables Bellec (2025).
[…]